OpenStudy (anonymous):

A website requires a four-digit numerical password in which the digits cannot repeat. How many possible passwords are there? 24 5,040 151,200 3,628,800

OpenStudy (anonymous):

10!=3628800

OpenStudy (ash2326):

We have 4 positions \[\large ----\] We have digits 0-9, total 10 digits

OpenStudy (anonymous):

10x9x8x7x6x5x4x3x2x1

OpenStudy (ash2326):

First position can be filled in 10 ways after choosing one no., we have 9 digits for the second place and similarly 8 for the third and 7 for the fourth \[\huge10 \times 9 \times 8 \times 7=5040\]

OpenStudy (anonymous):

oh yea

OpenStudy (ash2326):

@Lunalovegood do you understand?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

Good:)